OpenStudy (anonymous):

Solve for x: 8^{2x+3} = 32^{x+2}

OpenStudy (anonymous):

use natural logs. drop the power and solve like a normal equation

OpenStudy (jdoe0001):

\(\bf 8^{2x+3} = 32^{x+2}\qquad \qquad {\color{blue}{ 8\implies 2^3\qquad 32\implies 2^5}}\qquad thus \\ \quad \\ 8^{2x+3} = 32^{x+2}\implies ({\color{blue}{ 2^3}})^{2x+3}=({\color{blue}{ 2^5}})^{x+2}\) what do you think?

OpenStudy (anonymous):

there you go!

OpenStudy (anonymous):

Thanks drag0night

OpenStudy (anonymous):

No problem!